In an experiment, to determine the percentage purity of the hydrochloric acid, 25cm3 of the acid solution made by dissolving 25.5g of impure acid in a liter of solution was found to neutralize 25cm3 of 0.4M NaOH solution. Calculate
a. Concentration of the acid in
(i) g/dm3
(ii) mol/dm3
b. percentage purity of the acid
c. percentage impurity of the acid
d. What mass of impurities contained in 25.5g?
1
Expert's answer
2020-06-29T07:45:09-0400

V1S1=V2S2        Here, V1= 25 cm3, V2=25 cm3, S2= 0.4 M

S1 = (25×0.4) /25 M

    =0.4 M

Strength of the acid is 0.4 mol / dm3

m = n×M              Here, n= 0.4 mol / dm3, M= 36.46 g/mol

  = 0.4×36.46 g / dm3

   =14.584 g / dm3


a)

i) Concentration of the acid is 14.584 g / dm3 (Answer)

ii) Concentration of the acid is 0.4 mol / dm3 (Answer)

b)

percentage purity of the acid = (14.584/25.5) ×100

                                             = 57.19 percent (Answer)

c)

percentage impurity of the acid = 100-57.19 = 42.81 percent (Answer)


d)

mass of impurities contained in 25.5g = 25.5-14.584 g

                                                                  =10.916 g (Answer)